Could you please calculate the NPV. I always get the wrong answer. Tom now considers getting a two-year graduate degree - Tuition is $10,000 per year payable at the beginning of each year - To
Answer #1

B2 X fic end of year end of year cash flow present value factor @ 4% | present value of cash flow $ 10,000.00 1.000 $ 10,000.

for formulas and calculations, refer to the image below -

B2 X face end of year B wMDN 00 end of yearl cash flow present value factor @ 4% -10000 -10000 =D3/1.04 10000 =D4/1.04 104000

In case you have any query, kindly ask in comments.